OpenStudy (anonymous):

solve by elimination x-y-2z=4 -x+2y+z=1 -x+y-3z=11 please help i forgot how to do it

OpenStudy (anonymous):

Give me a sec, its a long process lol

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

do you HAVE to use elimination?? lol

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Damn.

OpenStudy (anonymous):

can you do this one ? a+b+c=-3 3b-c=4 2a-b-2c=-5

OpenStudy (anonymous):

OH! I got it, I was being stupid for a sec...

OpenStudy (anonymous):

so can you explain it to me?

OpenStudy (anonymous):

http://www.youtube.com/watch?v=aPKz7011lAA

OpenStudy (anonymous):

what did you multiply it by?

OpenStudy (anonymous):

watch the video.. I can't explain it.. lol sorry man I thought I knew